# This version: August 29, 2025 )

### v3

@@ A.3 Sharp Bounds on Treatment Effects
In this section, we utilize Artstein’s inequality (Theorem A.1 from Molinari (2020)) to characterize sharp bounds on the joint distribution of potential outcomes and true local preferences. That set of distributions produces sharp bounds on the averages of the treatment effects Y(j)Y(k)Y(j)-Y(k) conditional on $Q_{j}=(j,k)$ and $S_{j}=c_{j}$ for any comparable pair $(j,k)\in\mathcal{P}$. Recall that we denote the space of possible outcomes $Y$ as $\mathcal{Y}$. Let $2^{\mathcal{J}^{0}\times\mathcal{J}^{0}}$ denote the power set of $\mathcal{J}^{0}\times\mathcal{J}^{0}$} and $2^{\mathcal{Y}}$ denote the power set of $\mathcal{Y}$}.
@@
Proof of Theorem 1: The random set $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ is a measurable map from $\Omega$ to $\mathcal{Y}\times\mathcal{J}^{0}\times\mathcal{J}^{0}$, so it is compact valued. Following Definition A.1 from Molinari (2020), we say that $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ is a random closed set because, for every compact set $\mathcal{K}\in\mathbb{R}^{3}$, the set $\{\omega\in\Omega:(\{Y(\omega)\}\times\bm{\mathrm{Q}}_{j}(\omega))\cap\mathcal{K}\neq\emptyset\}$ is a measurable event. By Assumption 6(i), the random vector $(Y,Q_{j})$ and the random set $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ are measurable maps on the same probability space, and $\mathbb{P}\left[(Y,Q_{j})\in(\{Y\}\times\bm{\mathrm{Q}}_{j})|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$.
Artstein’s inequality (Theorem A.1 from Molinari (2020)) characterizes the sharp set of all possible probability distributions for $(Y,Q_{j})$ that are consistent with our observation of $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ and the fact that $\mathbb{P}[(Y,Q_{j})\in(\{Y\}\times\bm{\mathrm{Q}}_{j})|S_{j}=s]=1$. For any $s\in\mathcal{S}_{j}$, the inequality says that
@@
Lemma 1 of Chesher and Rosen (2017) applied to our case shows that the inequalities (A-3) and (A-5) are equivalent to:
@@
In case of continuous $Y$, it is impossible to directly evaluate all inequalities of Theorem 1 because there are uncountably many sets $A\in 2^{\mathcal{Y}}$. This is one of the drawbacks of the Artstein inequality approach, which has been extensively discussed by Beresteanu et al. (2012). In case $Y$ takes finitely many values, e.g., when $Y$ is binary, the number of such inequalities is feasible to evaluate because $2^{\mathcal{Y}}\times\bm{\mathrm{\Lambda}}^{\cup+}_{j}$ (or $2^{\mathcal{Y}}\times\bm{\mathrm{\Lambda}}^{\cup-}_{j}$) has finitely many elements. In fact, the number of inequalities is slightly higher than the number of inequalities in Proposition 4 of the main text, which we utilize to compute lower bounds on $\delta_{j,k}^{+}$ and $\delta_{j,k}^{-}$.
A.4 Agents that Maximize Expected Utility
This section micro-founds the strategic reporting behavior of agents by describing in detail the problem they have to solve. We build on Agarwal and Somaini (2018) and Fack et al. (2019), who propose expected-utility maximizing agents that calculate probabilities of admission for each possible choice of rank-ordered list (ROL or $P$). In their settings, each agent knows her type, the mechanism, the distribution of other agents’ types, but does not know the individual types of other agents and the $P$’s they decide to submit. Thus, the uncertainty about an agent’s match comes from this lack of information about other agents.
@@ A.4 Agents that Maximize Expected Utility
A.5 Issues When Controlling for $P$
@@ B.3 Proof of Proposition 2
Case 1.2.1: If $a\neq b$, we have $aQb$. We can always find a $Q$ such that $dQb$ and the second coordinate of $Q_{j}$ equals $d$; and we can always find another $Q$ such that $bQd$ and the second coordinate of $Q_{j}$ equals $b$. Therefore, $\bm{\mathrm{Q}}_{j}=\{P_{j}\}\cup\{(a\times N_{j}^{-})\}$.
@@ B.5 Proof of Corollary 1
$b\bar{P}d$ implies that $b\bar{Q}d$. Given that $dQe$, we have $bQe$. In case $e\in N_{j}^{-}$, it is no longer true that we can construct $Q$ such that $fQl$ for any $f\in N_{j}^{-}$. We can do so only for $f\neq e$. Therefore, $Q_{j}=(a,e)$ does not belong to $\bm{\mathrm{Q}}_{j}$.
@@
Sharpness follows because we remove all $Q$s that violate the implication of Proposition 3.
$\square$
@@ B.6 Proof of Proposition 4
The random set $\bm{\mathrm{Q}}_{j}$ is a measurable map from $\Omega$ to $\mathcal{J}^{0}\times\mathcal{J}^{0}$, so it is compact valued. Following Definition A.1 by Molinari (2020), we say that $\bm{\mathrm{Q}}_{j}$ is a random closed set because for every compact set $\mathcal{K}\in\mathbb{R}^{2}$, the set $\{\omega\in\Omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\mathcal{K}\neq\emptyset\}$ is a measurable event. By Assumption 6(i), the random vector $Q_{j}$ and the random set $\bm{\mathrm{Q}}_{j}$ are measurable maps on the same probability space, and $\mathbb{P}\left[Q_{j}\in\bm{\mathrm{Q}}_{j}|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$.
Artstein’s inequality (Theorem A.1 by Molinari (2020)) characterizes all possible probability mass functions $\mathbb{P}\left[Q_{j}=(a,b)|S_{j}=s\right]$ over $(a,b)\in\mathcal{J}^{0}\times\mathcal{J}^{0}$ for any $Q_{j}$ such that $\mathbb{P}\left[Q_{j}\in\bm{\mathrm{Q}}_{j}|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$. Since $\bm{\mathrm{Q}}_{j}$ is sharp (Assumption 6(ii)), the Artstein’s inequality yields the sharp set of all probability mass functions $\mathbb{P}\left[Q_{j}=(a,b)|S_{j}=s\right]$. For any $s\in\mathcal{S}_{j}$, the inequality says that
@@
Lemma 1 of Chesher and Rosen (2017) applied to our case shows that the inequalities (B-5) are equivalent to:
@@ Appendix C Empirical Appendix
SPO WPO WPO WPO
UMAS UMAS
fields
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 155 108 140 181 1001 370 1007 736
Share listing <<8 choices .38 .38 .8 .96 .9 .97 .9 .91
$\underline{\delta}_{j,k}$ 1 .43 .92 .83 .89 .28 .89 .04
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 104 77 201 34 239 89 241 334
Share listing <<8 choices .75 .18 .89 .91 .89 .94 .89 .81
$\underline{\delta}_{j,k}$ .84 .58 .43 1 .36 .51 .36 .13
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 127 63 111 111 936 250 957 316
Share listing <<8 choices .14 .41 .72 .86 .89 .9 .88 .88
$\underline{\delta}_{j,k}$ .71 .42 .23 .68 .18 .29 .16 .07
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 118 28 307 30 544 94 622 151
Share listing <<8 choices .33 .53 .87 .83 .89 .93 .87 .91
$\underline{\delta}_{j,k}$ .39 1 .15 1 .08 .34 .07 .2
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 50 86 36 235 36 235 148 352
Share listing <<8 choices .51 .68 .72 .92 .72 .92 .83 .92
$\underline{\delta}_{j,k}$ 1 .71 1 .15 1 .15 .24 .09
@@
SPO WPO WPO WPO
UMAS UMAS
fields
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 382 390 537 427 537 427 537 427
Share listing <<8 choices .88 .92 .91 .93 .91 .93 .91 .93
$\underline{\delta}_{j,k}$ 1 .92 .67 .31 .67 .31 .67 .31
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 76 23 255 182 255 182 255 372
Share listing <<8 choices .57 .47 .87 .96 .87 .96 .87 .96
$\underline{\delta}_{j,k}$ .71 .87 .21 .11 .21 .11 .21 .05
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 61 14 123 38 123 38 123 352
Share listing <<8 choices .7 .07 .85 .97 .85 .97 .85 .96
$\underline{\delta}_{j,k}$ .86 1 .43 .51 .43 .51 .43 .05
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 70 32 72 58 77 147 77 400
Share listing <<8 choices .89 .18 .9 .94 .9 .96 .9 .93
$\underline{\delta}_{j,k}$ 1 .81 .97 .45 .9 .17 .9 .06
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 72 41 93 92 967 175 967 395
Share listing <<8 choices .51 .31 .86 .92 .96 .93 .96 .92
$\underline{\delta}_{j,k}$ .58 .38 .45 .17 .04 .08 .04 .03

### v4

@@ A.3 Sharp Bounds on Treatment Effects
In this section, we utilize Artstein’s inequality (Theorem A.1 from molinari2020) to characterize sharp bounds on the joint distribution of potential outcomes and true local preferences. That set of distributions produces sharp bounds on the averages of the treatment effects Y(j)Y(k)Y(j)-Y(k) conditional on $Q_{j}=(j,k)$ and $S_{j}=c_{j}$ for any comparable pair $(j,k)\in\mathcal{P}$. Recall that we denote the space of possible outcomes $Y$ as $\mathcal{Y}$. Let $2^{\mathcal{J}^{0}\times\mathcal{J}^{0}}$ denote the power set of $\mathcal{J}^{0}\times\mathcal{J}^{0}$} and $2^{\mathcal{Y}}$ denote the power set of $\mathcal{Y}$}.
@@
Proof of Theorem 1: The random set $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ is a measurable map from $\Omega$ to $\mathcal{Y}\times\mathcal{J}^{0}\times\mathcal{J}^{0}$, so it is compact valued. Following Definition A.1 from molinari2020, we say that $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ is a random closed set because, for every compact set $\mathcal{K}\in\mathbb{R}^{3}$, the set $\{\omega\in\Omega:(\{Y(\omega)\}\times\bm{\mathrm{Q}}_{j}(\omega))\cap\mathcal{K}\neq\emptyset\}$ is a measurable event. By Assumption 6(i), the random vector $(Y,Q_{j})$ and the random set $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ are measurable maps on the same probability space, and $\mathbb{P}\left[(Y,Q_{j})\in(\{Y\}\times\bm{\mathrm{Q}}_{j})|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$.
Artstein’s inequality (Theorem A.1 from molinari2020) characterizes the sharp set of all possible probability distributions for $(Y,Q_{j})$ that are consistent with our observation of $(\{Y\}\times\bm{\mathrm{Q}}_{j})$ and the fact that $\mathbb{P}[(Y,Q_{j})\in(\{Y\}\times\bm{\mathrm{Q}}_{j})|S_{j}=s]=1$. For any $s\in\mathcal{S}_{j}$, the inequality says that
@@
Lemma 1 of chesher_ecta2017 applied to our case shows that the inequalities (A-3) and (A-5) are equivalent to:
@@
In case of continuous $Y$, it is impossible to directly evaluate all inequalities of Theorem 1 because there are uncountably many sets $A\in 2^{\mathcal{Y}}$. This is one of the drawbacks of the Artstein inequality approach, which has been extensively discussed by beresteanu_joe2012. In case $Y$ takes finitely many values, e.g., when $Y$ is binary, the number of such inequalities is feasible to evaluate because $2^{\mathcal{Y}}\times\bm{\mathrm{\Lambda}}^{\cup+}_{j}$ (or $2^{\mathcal{Y}}\times\bm{\mathrm{\Lambda}}^{\cup-}_{j}$) has finitely many elements. In fact, the number of inequalities is slightly higher than the number of inequalities in Proposition 4 of the main text, which we utilize to compute lower bounds on $\delta_{j,k}^{+}$ and $\delta_{j,k}^{-}$.
A.4 Falsification Test
It is useful to characterize a falsification test based on the assumptions that we employ to construct $\bm{\mathrm{Q}}_{j}$. Such a test can rely on the fact that the right-hand sides of the inequalities in Proposition 4 must provide a lower bound for a probability mass function if the model assumptions are correct. For any partition $\mathcal{A}$ of $\mathcal{J}^{0}\times\mathcal{J}^{0}$, we must have $\sum_{A\in\mathcal{A}}\mathbb{P}\left[Q_{j}\in A|S_{j}=c_{j}\right]=1$ for any given distribution in the sharp set of Proposition 4. Thus, the same sum applied to the right-hand sides of the inequalities above must be less than or equal to one.
Corollary 2 (Model’s Falsification Test).
A.5 Agents that Maximize Expected Utility
This section micro-founds the strategic reporting behavior of agents by describing in detail the problem they have to solve. We build on agarwal2018 and fack2019, who propose expected-utility maximizing agents that calculate probabilities of admission for each possible choice of rank-ordered list (ROL or $P$). In their settings, each agent knows her type, the mechanism, the distribution of other agents’ types, but does not know the individual types of other agents and the $P$’s they decide to submit. Thus, the uncertainty about an agent’s match comes from this lack of information about other agents.
@@ A.5 Agents that Maximize Expected Utility
A.6 Issues When Controlling for $P$
@@ B.3 Proof of Proposition 2
Case 1.2.1: If $a\neq b$, we have $aQb$. We can always find a $Q$ such that $dQb$ and the second coordinate of $Q_{j}$ equals $d$; and we can always find another $Q$ such that $bQd$ and the second coordinate of $Q_{j}$ equals $b$. Therefore, $\bm{\mathrm{Q}}_{j}=\{P_{j}\}\cup\{(\{a\}\times N_{j}^{-})\}$.
@@ B.5 Proof of Corollary 1
$b\bar{P}d$ implies that $b\bar{Q}d$. Given that $dQe$, we have $bQe$. In case $e\in N_{j}^{-}$, it is no longer true that we can construct $Q$ such that $fQb$ for any $f\in N_{j}^{-}$. We can do so only for $f\neq e$. Therefore, $Q_{j}=(a,e)$ does not belong to $\bm{\mathrm{Q}}_{j}$.
@@
Sharpness follows because we remove all $Q$s that violate the implication of Prop. 3. $\square$
@@ B.6 Proof of Proposition 4
The random set $\bm{\mathrm{Q}}_{j}$ is a measurable map from $\Omega$ to $\mathcal{J}^{0}\times\mathcal{J}^{0}$, so it is compact valued. Following Definition A.1 by molinari2020, we say that $\bm{\mathrm{Q}}_{j}$ is a random closed set because for every compact set $\mathcal{K}\in\mathbb{R}^{2}$, the set $\{\omega\in\Omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\mathcal{K}\neq\emptyset\}$ is a measurable event. By Assumption 6(i), the random vector $Q_{j}$ and the random set $\bm{\mathrm{Q}}_{j}$ are measurable maps on the same probability space, and $\mathbb{P}\left[Q_{j}\in\bm{\mathrm{Q}}_{j}|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$.
Artstein’s inequality (Theorem A.1 by molinari2020) characterizes all possible probability mass functions $\mathbb{P}\left[Q_{j}=(a,b)|S_{j}=s\right]$ over $(a,b)\in\mathcal{J}^{0}\times\mathcal{J}^{0}$ for any $Q_{j}$ such that $\mathbb{P}\left[Q_{j}\in\bm{\mathrm{Q}}_{j}|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$. Since $\bm{\mathrm{Q}}_{j}$ is sharp (Assumption 6(ii)), the Artstein’s inequality yields the sharp set of all probability mass functions $\mathbb{P}\left[Q_{j}=(a,b)|S_{j}=s\right]$. For any $s\in\mathcal{S}_{j}$, the inequality says that
@@
Lemma 1 of chesher_ecta2017 applied to our case shows that the inequalities (B-5) are equivalent to:
@@ Appendix C Empirical Appendix
SPO SPO WPO WPO
UMAS UMAS UMAS UMAS
fields fields
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 87 48 153 51 140 181 1001 370
Share listing <<8 choices .67 .87 .38 .82 .8 .96 .9 .97
$\underline{\delta}_{j,k}$ 1 1 1 1 .92 .83 .89 .28
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 101 17 104 19 201 34 239 89
Share listing <<8 choices .78 .82 .75 .73 .89 .91 .89 .94
$\underline{\delta}_{j,k}$ .87 1 .84 1 .43 1 .36 .51
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 49 41 121 49 111 111 936 250
Share listing <<8 choices .38 .63 .15 .53 .72 .86 .89 .9
$\underline{\delta}_{j,k}$ .86 1 .75 1 .23 .68 .18 .29
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 78 20 99 21 307 30 544 94
Share listing <<8 choices .51 .75 .4 .71 .87 .83 .89 .93
$\underline{\delta}_{j,k}$ .6 1 .47 1 .15 1 .08 .34
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 36 77 36 77 36 235 36 235
Share listing <<8 choices .72 .76 .72 .76 .72 .92 .72 .92
$\underline{\delta}_{j,k}$ 1 .8 1 .8 1 .15 1 .15
@@
SPO SPO WPO WPO
UMAS UMAS UMAS UMAS
fields fields
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 382 390 382 390 537 427 537 427
Share listing <<8 choices .88 .92 .88 .92 .91 .93 .91 .93
$\underline{\delta}_{j,k}$ 1 .92 1 .92 .67 .31 .67 .31
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 76 17 76 17 255 182 255 182
Share listing <<8 choices .57 .64 .57 .64 .87 .96 .87 .96
$\underline{\delta}_{j,k}$ .71 1 .71 1 .21 .11 .21 .11
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 61 2 61 2 123 38 123 38
Share listing <<8 choices .7 .5 .7 .5 .85 .97 .85 .97
$\underline{\delta}_{j,k}$ .86 1 .86 1 .43 .51 .43 .51
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 70 9 70 11 72 58 77 147
Share listing <<8 choices .89 .66 .89 .54 .9 .94 .9 .96
$\underline{\delta}_{j,k}$ 1 1 1 1 .97 .45 .9 .17
@@
Obs. with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 50 20 72 25 93 92 967 175
Share listing <<8 choices .74 .64 .51 .51 .86 .92 .96 .93
$\underline{\delta}_{j,k}$ .83 .78 .58 .62 .45 .17 .04 .08